# On the contraction properties of Sinkhorn semigroups

### v2

@@ 1.3 Contributions
These powerful off-the-shelf semigroup techniques are mainly based on two conditions: 1) the existence of a Lyapunov function satisfying a drift-type condition; 2) a local minorization condition on the sub-level sets of the Lyapunov function.
@@
Such drift-minorization approaches are well-established techniques to localize the minorization conditions on the compact sub-level sets of the Lyapunov function. In this context, Markov operators act on weighted Banach spaces and operator norms coincide with the contraction coefficients of the Markov integral operator. For a more detailed discussion on this contraction coefficients we refer to [dpa, horton-dp, penev].
@@
This Lyapunov approach and the operator-theoretic framework presented in this article strongly differ from the approaches discussed earlier and taken in the series of recent articles [chiarini, durmus, del2025stability, eckstein2025hilbert].
Our main contributions can be summarized in the following way:
@@
  • In the context of bounded transition potentials, we provide novel strict contraction estimates with respect to general $\Phi$-divergence criteria, including $\mathbb{L}_{p}$-norms, Havrda-Charvat entropies, $\alpha$-divergences, Jensen-Shannon divergence, squared Hellinger integrals and Kakutani-Hellinger, as well as Jeffreys and Rényi divergences [csiszar, dm-03, majernik, nielsen] (see, for instance, Theorem 2.2 and Section 4.1).
@@
  • In the context of locally bounded cost functions and marginal potentials with compact sub-level sets, we develop new weighted total variation norm estimates, including Kantorovich and Wasserstein-type distances (see Section 4.4 and Section 4.5) as well as relative entropy bridge estimates (in Section 4.6).
    We also emphasize that the choice of the Wasserstein semi-metric and the Lyapunov functions depends on the marginal target measures (see Theorem 2.3 and Section 4.5).
  • We develop general Lyapunov contractions principles under practical and minimal regularity conditions that allow to provide quantitative exponential stability estimates for a large class of models.
    We apply this novel framework in a variety of situations, ranging from polynomial growth potentials and heavy tailed marginals on general normed spaces to more sophisticated boundary state space models, including semi-circle transitions, Beta, Weibull, exponential marginals as well as semi-compact models (e.g., semi-discrete marginal models). Last but not least, our approach also allows to consider statistical finite mixture of the above models including kernel-type density estimators of complex data distributions arising in generative models.
To the best of our knowledge, the stability results discussed above are the first results of this type in the literature on entropic optimal transport and Sinkhorn semigroups.
1.4 Outline of the article
After presenting some basic notation in Section 1.5, we begin with a brief review on Schrödinger bridge problems and Sinkhorn semigroups in Section 2 (and we refer to [adm-24] for a more detailed discussion on these models within our framework).
Section 2.3, presents a brief summary of the main stability theorems of the article. All the methodologies and regularity conditions discussed in the article are illustrated in a variety of examples in Section 5. To motivate our results, however, Section 2.4 already includes the discussion of some application models. We also give some detailed comments on our regularity conditions and propose an avenue of open research projects.
Some background on Lyapunov approaches and weighted-norm contraction properties of Markov integral operators is provided in Section 3, with additional details in Appendix C. Section 3.1 recalls some basics facts on the Hilbert projective metric, operator norms, entropy and related $\phi$-divergences. The operator contraction coefficients associated with these criteria are introduced in Section 3.2.
With these notations at hand, Section 4 shows how to adapt this operator-theoretic framework to the convergence analysis of Sinkhorn semigroups.
Strong mixing Sinkhorn semigroups are discussed in Section 4.1 in terms of Dobrushin contraction coefficients. In this context, we show that Sinkhorn semigroups contract for any $\phi$-divergence criteria, including the relative entropy, the total variation distance, as well as $\mathbb{L}_{p}$-norms, Havrda-Charvat entropies, $\alpha$-divergences, Jensen-Shannon divergence, squared Hellinger integrals and Kakutani-Hellinger, Jeffreys and Rényi divergences.
Section 4.2 presents some general drift-minorization conditions that enable us to estimate the contraction coefficients of Sinkhorn semigroups. The main contraction theorems introduced in this article are presented in Section 4.3. Section 4.4 presents a simple and new regularity condition that allows to provide quantitative exponential stability estimates for a large class of models. Section 4.5 presents quantitative exponential decays w.r.t. Wasserstein semi-metrics and Section 4.6 is dedicated to relative entropy decays and the convergence of Sinkhorn bridges.
Section 5 is dedicated to a variety of illustrations. In particular, polynomial growth models are discussed in Section 5.1, Section 5.2 is dedicated to heavy tailed models, and in Section 5.3 we study bounded Lipschitz domains. Beta and Weibull distributions are discussed, respectively, in Section 5.4 and Section 5.5. An explicit description of the key regularity conditions and results in the context of multivariate linear Gaussian distributions is provided in Section 5.8. Finally, Section 5.9 is dedicated to statistical finite (non necessarily Gaussian) mixture models.
Detailed proofs and some extended analyses can be found in the appendices to this manuscript. Appendix A is devoted to the Gaussian Sinkhorn algorithm, Appendices B and C contain material related to Hilbert metric and weighted norm contractions, respectively. Appendix D is devoted to Schrödinger potential functions. Finally, some additional technical proofs provided for completeness in Appendix E.
1.5 Some basic notation
@@
We let ${\cal B}(\mathbb{X})$ be the algebra of locally bounded measurable functions on a locally compact Polish space $\mathbb{X}$. We denote by ${\cal B}_{b}(\mathbb{X})\subset{\cal B}(\mathbb{X})$ the sub-algebra of bounded measurable functions $f$ on $\mathbb{X}$ endowed with the supremum norm $\|f\|:=\sup_{x\in\mathbb{X}}|f(x)|$. For a given uniformly positive function $\varphi\in{\cal B}(\mathbb{X})$, we let ${\cal B}_{\varphi}(\mathbb{X})\subset{\cal B}(\mathbb{X})$ be the Banach space of functions $f\in{\cal B}(\mathbb{X})$ equipped with the norm $\|f\|_{\varphi}:=\|f/\varphi\|$.
@@
Let ${\cal M}(\mathbb{X})$ be the set of locally bounded signed measures on $\mathbb{X}$, and let ${\cal M}_{b}(\mathbb{X})\subset{\cal M}(\mathbb{X})$ the subset of bounded measures. We denote the convex subset of probability measures on $\mathbb{X}$ as ${\cal M}_{1}(\mathbb{X})\subset{\cal M}_{b}(\mathbb{X})$.
The Lebesgue integral of an integrable function $f\in{\cal B}(\mathbb{X})$ w.r.t. $\nu\in{\cal M}(\mathbb{X})$ is denoted by
@@
Consider a bounded integral operator ${\cal K}(x,dy)$ from $\mathbb{X}$ into $\mathbb{Y}$, a measure $\mu\in{\cal M}(\mathbb{X})$ and a function $f\in{\cal B}(\mathbb{Y})$. We denote by $\mu{\cal K}\in{\cal M}(\mathbb{Y})$ and ${\cal K}(f)\in{\cal B}(\mathbb{X})$ the measure and the function defined by
$(\mu{\cal K})(dy):=\int~\mu(dx){\cal K}(x,dy)\quad\mbox{\rm and}\quad{\cal K}(f)(x):=\int{\cal K}(x,dy)f(y).$ (6)
For indicator functions $f=1_{A}$ sometimes we write ${\cal K}(x,A)$ instead of ${\cal K}(1_{A})(x)$. For bounded integral operator ${\cal K}_{1}(x,dy)$ from $\mathbb{X}$ into $\mathbb{Y}$ and ${\cal K}_{2}(y,dx)$ from $\mathbb{Y}$ into $\mathbb{X}$, we also denote by $({\cal K}_{1}{\cal K}_{2})$ the bounded integral operator from $\mathbb{X}$ into $\mathbb{X}$ defined by the integral composition
$({\cal K}_{1}{\cal K}_{2})(x_{1},dx_{2}):=\int{\cal K}_{1}(x_{1},dy){\cal K}_{2}(y,dx_{2}).$
Given a Markov transition ${\cal K}(x,dy)$ from $\mathbb{X}$ into $\mathbb{Y}$ and a measure $\mu\in{\cal M}(\mathbb{X})$ we set
$(\mu\times{\cal K})(d(x,y)):=\mu(dx)\,{\cal K}(x,dy)\quad\mbox{\rm and}\quad(\mu\times{\cal K})^{\flat}(d(y,x)):=\mu(dx)\,{\cal K}(x,dy).$
A measure $\nu_{1}\in{\cal M}(\mathbb{X})$ is said to be absolutely continuous with respect to another measure $\nu_{2}\in{\cal M}(\mathbb{X})$ and we write $\nu_{1}\ll\nu_{2}$ if $\nu_{2}(A)=0$ implies $\nu_{1}(A)=0$ for any $A\in\mathbb{X}$. Whenever $\nu_{1}\ll\nu_{2}$, we denote by ${d\nu_{1}}/{d\nu_{2}}$ the Radon-Nikodym derivative of $\nu_{1}$ w.r.t. $\nu_{2}$. We also consider the partial order, denoted $\nu_{1}\leq\nu_{2}$, when $\nu_{1}(A)\leq\nu_{2}(A)$ for any measurable subset $A\subset\mathbb{X}$. The measures are said comparable when we have $c~\nu_{1}\leq\nu_{2}\leq c^{-1}~\nu_{1}$ for some scalar constant $c>0$.
@@
We consider the weighted total variation norm associated with some positive function $\varphi\in{\cal B}(\mathbb{X})$ defined as
@@
where $\Pi(\mu_{1},\mu_{2})$ stands for the set of coupling probability measures $\cal P$} on the product space $(\mathbb{X}\times\mathbb{X})$, with prescribed first and second coordinate marginals $(\mu_{1},\mu_{2})$ such that $\mu_{i}(\varphi)<\infty$. For a given semi-metric $\varrho(x_{1},x_{2})$ on $\mathbb{X}$ and $p\geq 1$ we also consider the Wasserstein semi-metric $\mathbb{W}_{p,\varrho}$ defined by
$\mathbb{W}_{p,\varrho}(\mu_{1},\mu_{2}):=\inf_{{\cal P}\in\Pi(\mu_{1},\mu_{2})}\left(\int~\varrho(x_{1},x_{2})^{p}~{\cal P}(d(x_{1},x_{2}))\right)^{1/p}.$ (8)
@@
$\varphi(x_{1})+\varphi(x_{2})\geq\varrho(x_{1},x_{2})\quad\mbox{\rm and}\quad\varphi(x)\geq c>0$ (9)
@@
The choice of the semi-metric $\varrho$ in (10) is not unique. Indeed, developing the exponential function for any $q\geq 1$ we also have the estimate
$\varphi(x_{1})+\varphi(x_{2})\geq\varrho_{p,q}(x_{1},x_{2}):=c_{p,q}~\|x_{1}-x_{2}\|^{pq}\quad\mbox{\rm with}\quad c_{p,q}:=2c~\frac{1}{q!}\left(\frac{b}{2^{1+(p-1)_{+}}}\right)^{q},$ (11)
@@
$\begin{array}[]{cccl}\Phi:&\mathbb{R}_{+}^{2}&\mapsto&\mathbb{R}\cup\{+\infty\}\\ &(u,v)&\leadsto&\Phi(u,v)\quad\mbox{\rm s.t. $\forall a\in\mathbb{R}_{+}$~~~ $\Phi(au,av)=a~\Phi(u,v)$ and $\Phi(1,1)=0$.}\end{array}$ (12)
The $\Phi$-entropy between two probability measures $\nu_{1},\nu_{2}\in{\cal M}_{1}(S)$ is defined for any dominating measure $\gamma$ (such that $\gamma\gg\nu_{1}$ and $\gamma\gg\nu_{2}$) by the formula
$D_{\Phi}(\nu_{1},\nu_{2}):=\int~\Phi\left(\frac{d\nu_{1}}{d\gamma}(x),\frac{d\nu_{2}}{d\gamma}(x)\right)~\gamma(dx).$ (13)
By homogeneity arguments, the definition (13) does not depend on the choice of the dominating measure $\gamma$. Note that
$\text{if}\quad\Phi(u,v)=\Phi_{0}(u,v):=\frac{1}{2}~|u-v|\quad\text{then}\quad D_{\Phi_{0}}(\nu_{1},\nu_{2})=\|\nu_{1}-\nu_{2}\|_{\sf tv}.$ (14)
@@
2 Statement of main results
2.1 Sinkhorn bridges
The Schrödinger bridge distribution $\mathbb{P}$ defined in (4) can rarely be computed analytically. However, solutions can be approximated efficiently using the Sinkhorn algorithm, also referred to as the iterative proportional fitting procedure [cuturi, sinkhorn-2, sinkhorn-3]. This procedure is defined sequentially for any $n\geq 0$ by a collection of probability distributions (${\cal P}_{n}$) and transition operators (${\cal S}_{n}$) given by
${\cal P}_{2n}=\lambda_{U}\times{\cal S}_{2n}\quad\mbox{\rm and}\quad{\cal P}_{2n+1}=(\nu_{V}\times{\cal S}_{2n+1})^{\flat}$ (15)
starting from ${\cal P}_{0}$ with ${\cal S}_{0}={\cal Q}$, where, for a measure $\mu(dy)$ and a transition operator ${\cal S}(y,dx)$, the notation $(\mu\times{\cal S})^{\flat}$ denotes the ‘reverse’ measure
$(\mu\times{\cal S})^{\flat}(d(x,y)):=\mu(dy){\cal S}(y,dx).$
The Markov transitions ${\cal S}_{n}$ in (15) are defined by the Sinkhorn iterations
$\left\{\begin{array}[]{l}(\pi_{2n}\times{\cal S}_{2n+1})^{\flat}=\lambda_{U}\times{\cal S}_{2n}\quad\mbox{\rm and}\quad\pi_{2n+1}\times{\cal S}_{2(n+1)}=(\nu_{V}\times{\cal S}_{2n+1})^{\flat}\\ \\ \mbox{\rm with the distributions}\quad\pi_{2n}:=\lambda_{U}{\cal S}_{2n}\quad\mbox{\rm and}\quad\pi_{2n+1}:=\nu_{V}{\cal S}_{2n+1}.\end{array}\right$. (16)
In the dual formulation, the bridge distributions ${\cal P}_{n}$ are often written as
${\cal P}_{n}(d(x,y))=e^{-W(x,y)}~\lambda_{U_{n}}(dx)~\nu_{V_{n}}(dy),$ (17)

### v3

@@ 1.3. Contributions
These powerful off-the-shelf semigroup techniques are mainly based on two conditions:
  1. (1)
    The existence of a Lyapunov function satisfying a drift-type condition.
  2. (2)
    A local minorization condition on the sub-level sets of the Lyapunov function.
@@
Such drift-minorization approaches are well-established techniques to localize the minorization conditions on the compact sub-level sets of the Lyapunov function. In this context, Markov operators act on weighted Banach spaces and operator norms coincide with the contraction coefficients of the Markov integral operator. For a more detailed discussion on this contraction coefficients we refer to [dpa, delgerber2025, horton-dp, penev].
@@
This Lyapunov approach and the operator-theoretic framework presented in this article strongly differ from the approaches discussed earlier and taken in the series of recent articles [chiarini, durmus, del2025stability, eckstein2025hilbert]. Our main contributions can be summarized in the following way:
@@
  • Strict contraction for bounded costs: For bounded transition potentials, we establish strict contraction inequalities for a broad class of $\phi$-divergences, including $\mathbb{L}_{p}$ norms, Havrda–Charvat and $\alpha$-divergences, Jensen–Shannon divergence, squared Hellinger and Kakutani–Hellinger distances, as well as Jeffreys and Rényi divergences (see Theorem 2.2 and Section 4.1).
@@
  • Weighted-norm stability for unbounded models: For locally bounded costs and marginal potentials with compact sub-level sets, we introduce weighted total variation and Kantorovich–Wasserstein contraction estimates, together with relative entropy bounds for Schrödinger bridges (Sections 4.4, 4.5, and 4.6). The choice of Lyapunov functions and Wasserstein metrics is explicitly adapted to the marginal distributions.
  • General Lyapunov contraction principles: We formulate practical drift–minorization conditions yielding quantitative exponential convergence rates for a large class of Sinkhorn semigroups. These results apply to models with polynomial growth, heavy-tailed marginals, boundary state spaces, semi-compact domains, and finite mixture models, including kernel-based density estimators used in generative modeling.
To the best of our knowledge, these results provide the first general exponential stability estimates for Sinkhorn semigroups with respect to weighted norms, $\phi$-divergences, and Wasserstein-type distances in non-necessarily compact and unbounded settings.
1.4. Outline of the article
After introducing notation in Section 1.5, Section 2 recalls the Schrödinger bridge problem and the associated Sinkhorn semigroups, emphasizing their interpretation as time-inhomogeneous nonlinear Markov evolutions (see also [adm-24]).
Section 2.3 presents a concise overview of the main stability and contraction results established in this work. Motivating applications are already discussed in Section 2.4, which highlights the relevance of our approach for non-compact state spaces, unbounded costs, and heavy-tailed marginals.
Sections 33.2 develop the operator-theoretic foundations underlying our analysis. We review Lyapunov drift–minorization techniques and weighted Banach space contractions for Markov integral operators (Section 3, with complements in Appendix C), recall the Hilbert projective metric and $\phi$-divergences (Section 3.1), and introduce the associated contraction coefficients (Section 3.2).
Building on this framework, Section 4 adapts weighted-norm and Lyapunov contraction techniques to Sinkhorn semigroups. Strong mixing regimes are analyzed in Section 4.1, where we establish strict contraction properties for a broad class of $\phi$-divergences, including relative entropy, total variation, $\mathbb{L}_{p}$ norms, Hellinger-type divergences, Jensen–Shannon, Jeffreys, and Rényi divergences. General drift–minorization conditions ensuring quantitative contraction bounds are developed in Section 4.2. The main exponential stability theorems are presented in Section 4.3. Section 4.4 introduces a simple and practically verifiable Lyapunov condition yielding exponential convergence on weighted Banach spaces. Extensions to Wasserstein-type semi-metrics are derived in Section 4.5, while Section 4.6 is devoted to relative-entropy decay and the convergence of Sinkhorn bridges.
Section 5 illustrates the scope of the theory through a wide range of examples, including polynomial growth models (Section 5.1), heavy-tailed marginals (Section 5.2), bounded Lipschitz domains (Section 5.3), Beta and Weibull models (Sections 5.4 and 5.5), and multivariate linear Gaussian models (Section 5.8). Section 5.9 extends the analysis to finite mixture models, including non-Gaussian and kernel-based constructions relevant to generative modeling.
Detailed proofs and complementary material are collected in the appendices. Appendix A focuses on Gaussian Sinkhorn algorithms, Appendices B and C provide additional results on Hilbert metrics and weighted-norm contractions, Appendix D reviews Schrödinger potentials, and Appendix E contains further technical arguments.
1.5. Some basic notation
@@
We let ${\mathcal{B}}(\mathbb{X})$ be the algebra of locally bounded measurable functions on a locally compact Polish space $\mathbb{X}$. We denote by ${\mathcal{B}}_{b}(\mathbb{X})\subset{\mathcal{B}}(\mathbb{X})$ the sub-algebra of bounded measurable functions $f$ on $\mathbb{X}$ endowed with the supremum norm $\|f\|:=\sup_{x\in\mathbb{X}}|f(x)|$. For a given uniformly positive function $\varphi\in{\mathcal{B}}(\mathbb{X})$, we let ${\mathcal{B}}_{\varphi}(\mathbb{X})\subset{\mathcal{B}}(\mathbb{X})$ be the Banach space of functions $f\in{\mathcal{B}}(\mathbb{X})$ equipped with the norm $\|f\|_{\varphi}:=\|f/\varphi\|$.
@@
Let ${\mathcal{M}}(\mathbb{X})$ be the set of locally bounded signed measures on $\mathbb{X}$, and let ${\mathcal{M}}_{b}(\mathbb{X})\subset{\mathcal{M}}(\mathbb{X})$ the subset of bounded measures. We denote the convex subset of probability measures on $\mathbb{X}$ as ${\mathcal{M}}_{1}(\mathbb{X})\subset{\mathcal{M}}_{b}(\mathbb{X})$.
The Lebesgue integral of an integrable function $f\in{\mathcal{B}}(\mathbb{X})$ w.r.t. $\nu\in{\mathcal{M}}(\mathbb{X})$ is denoted by
@@
Consider a bounded integral operator ${\mathcal{K}}(x,dy)$ from $\mathbb{X}$ into $\mathbb{Y}$, a measure $\mu\in{\mathcal{M}}(\mathbb{X})$ and a function $f\in{\mathcal{B}}(\mathbb{Y})$. We denote by $\mu{\mathcal{K}}\in{\mathcal{M}}(\mathbb{Y})$ and ${\mathcal{K}}(f)\in{\mathcal{B}}(\mathbb{X})$ the measure and the function defined by
(1.6) $(\mu{\mathcal{K}})(dy):=\int~\mu(dx){\mathcal{K}}(x,dy)\quad\mbox{\rm and}\quad{\mathcal{K}}(f)(x):=\int{\mathcal{K}}(x,dy)f(y).$
For indicator functions $f=1_{A}$ sometimes we write ${\mathcal{K}}(x,A)$ instead of ${\mathcal{K}}(1_{A})(x)$. For bounded integral operator ${\mathcal{K}}_{1}(x,dy)$ from $\mathbb{X}$ into $\mathbb{Y}$ and ${\mathcal{K}}_{2}(y,dx)$ from $\mathbb{Y}$ into $\mathbb{X}$, we also denote by $({\mathcal{K}}_{1}{\mathcal{K}}_{2})$ the bounded integral operator from $\mathbb{X}$ into $\mathbb{X}$ defined by the integral composition
$({\mathcal{K}}_{1}{\mathcal{K}}_{2})(x_{1},dx_{2}):=\int{\mathcal{K}}_{1}(x_{1},dy){\mathcal{K}}_{2}(y,dx_{2}).$
Given a Markov transition ${\mathcal{K}}(x,dy)$ from $\mathbb{X}$ into $\mathbb{Y}$ and a measure $\mu\in{\mathcal{M}}(\mathbb{X})$ we set
$(\mu\times{\mathcal{K}})(d(x,y)):=\mu(dx)\,{\mathcal{K}}(x,dy)\quad\mbox{\rm and}\quad(\mu\times{\mathcal{K}})^{\flat}(d(y,x)):=\mu(dx)\,{\mathcal{K}}(x,dy).$
A measure $\nu_{1}\in{\mathcal{M}}(\mathbb{X})$ is said to be absolutely continuous with respect to another measure $\nu_{2}\in{\mathcal{M}}(\mathbb{X})$ and we write $\nu_{1}\ll\nu_{2}$ if $\nu_{2}(A)=0$ implies $\nu_{1}(A)=0$ for any $A\in\mathbb{X}$. Whenever $\nu_{1}\ll\nu_{2}$, we denote by ${d\nu_{1}}/{d\nu_{2}}$ the Radon-Nikodym derivative of $\nu_{1}$ w.r.t. $\nu_{2}$. We also consider the partial order, denoted $\nu_{1}\leq\nu_{2}$, when $\nu_{1}(A)\leq\nu_{2}(A)$ for any measurable subset $A\subset\mathbb{X}$. The measures are said comparable when we have $c~\nu_{1}\leq\nu_{2}\leq c^{-1}~\nu_{1}$ for some scalar constant $c>0$.
@@
We consider the weighted total variation norm associated with some positive function $\varphi\in{\mathcal{B}}(\mathbb{X})$ defined as
@@
where $\Pi(\mu_{1},\mu_{2})$ stands for the set of coupling probability measures $\mathcal{P}$} on the product space $(\mathbb{X}\times\mathbb{X})$, with prescribed first and second coordinate marginals $(\mu_{1},\mu_{2})$ such that $\mu_{i}(\varphi)<\infty$. For a given semi-metric $\varrho(x_{1},x_{2})$ on $\mathbb{X}$ and $p\geq 1$ we also consider the Wasserstein semi-metric $\mathbb{W}_{p,\varrho}$ defined by
(1.8) $\mathbb{W}_{p,\varrho}(\mu_{1},\mu_{2}):=\inf_{{\mathcal{P}}\in\Pi(\mu_{1},\mu_{2})}\left(\int~\varrho(x_{1},x_{2})^{p}~{\mathcal{P}}(d(x_{1},x_{2}))\right)^{1/p}.$
@@
(1.9) $\varphi(x_{1})+\varphi(x_{2})\geq\varrho(x_{1},x_{2})\quad\mbox{\rm and}\quad\varphi(x)\geq c>0$
@@
The choice of the semi-metric $\varrho$ in (1.10) is not unique. Indeed, developing the exponential function for any $q\geq 1$ we also have the estimate
(1.11) $\varphi(x_{1})+\varphi(x_{2})\geq\varrho_{p,q}(x_{1},x_{2}):=c_{p,q}~\|x_{1}-x_{2}\|^{pq}\quad\mbox{\rm with}\quad c_{p,q}:=2c~\frac{1}{q!}\left(\frac{b}{2^{1+(p-1)_{+}}}\right)^{q},$
@@
(1.12) $\begin{array}[]{cccl}\Phi:&\mathbb{R}_{+}^{2}&\rightarrow&\mathbb{R}\cup\{+\infty\}\\ &(u,v)&\mapsto&\Phi(u,v)\quad\mbox{\rm s.t. $\forall a\in\mathbb{R}_{+}$~~~ $\Phi(au,av)=a~\Phi(u,v)$ and $\Phi(1,1)=0$.}\end{array}$
The $\Phi$-entropy between two probability measures $\nu_{1},\nu_{2}\in{\mathcal{M}}_{1}(S)$ is defined for any dominating measure $\gamma$ (such that $\gamma\gg\nu_{1}$ and $\gamma\gg\nu_{2}$) by the formula
(1.13) $D_{\Phi}(\nu_{1},\nu_{2}):=\int~\Phi\left(\frac{d\nu_{1}}{d\gamma}(x),\frac{d\nu_{2}}{d\gamma}(x)\right)~\gamma(dx).$
By homogeneity arguments, the definition (1.13) does not depend on the choice of the dominating measure $\gamma$. Note that
(1.14) $\text{if}\quad\Phi(u,v)=\Phi_{0}(u,v):=\frac{1}{2}~|u-v|\quad\text{then}\quad D_{\Phi_{0}}(\nu_{1},\nu_{2})=\|\nu_{1}-\nu_{2}\|_{\sf tv}.$
@@
2. Statement of main results
This section introduces the Sinkhorn bridge construction, the associated time-inhomogeneous Sinkhorn semigroups, and states our main exponential stability results. Detailed proofs and refinements are deferred to later sections.
2.1. Sinkhorn bridges
The Schrödinger bridge distribution $\mathbb{P}$ defined in (1.4) can rarely be computed analytically. However, solutions can be approximated efficiently using the Sinkhorn algorithm, also referred to as the iterative proportional fitting procedure [cuturi, sinkhorn-2, sinkhorn-3]. This procedure is defined sequentially for any $n\geq 0$ by a collection of probability distributions (${\mathcal{P}}_{n}$) and transition operators (${\mathcal{S}}_{n}$) given by
(2.1) ${\mathcal{P}}_{2n}=\lambda_{U}\times{\mathcal{S}}_{2n}\quad\mbox{\rm and}\quad{\mathcal{P}}_{2n+1}=(\nu_{V}\times{\mathcal{S}}_{2n+1})^{\flat}$
starting from ${\mathcal{P}}_{0}$ with ${\mathcal{S}}_{0}={\mathcal{Q}}$, where, for a measure $\mu(dy)$ and a transition operator ${\mathcal{S}}(y,dx)$, the notation $(\mu\times{\mathcal{S}})^{\flat}$ denotes the ‘reverse’ measure
$(\mu\times{\mathcal{S}})^{\flat}(d(x,y)):=\mu(dy){\mathcal{S}}(y,dx).$
The Markov transitions ${\mathcal{S}}_{n}$ in (2.1) are defined by the Sinkhorn iterations
(2.2) $\left\{\begin{array}[]{l}(\pi_{2n}\times{\mathcal{S}}_{2n+1})^{\flat}=\lambda_{U}\times{\mathcal{S}}_{2n}\quad\mbox{\rm and}\quad\pi_{2n+1}\times{\mathcal{S}}_{2(n+1)}=(\nu_{V}\times{\mathcal{S}}_{2n+1})^{\flat}\\ \\ \mbox{\rm with the distributions}\quad\pi_{2n}:=\lambda_{U}{\mathcal{S}}_{2n}\quad\mbox{\rm and}\quad\pi_{2n+1}:=\nu_{V}{\mathcal{S}}_{2n+1}.\end{array}\right$.
In the dual formulation, the bridge distributions ${\mathcal{P}}_{n}$ are often written as
(2.3) ${\mathcal{P}}_{n}(d(x,y))=e^{-W(x,y)}~\lambda_{U_{n}}(dx)~\nu_{V_{n}}(dy),$